{r = (2 = (Z -t)y,-3<t< 3; y(1) = 2 Find the Peano range of the Cauchy problem.

Z=38
Answer #1

y = (2-t) y y (1)= 2 yl 1) = 38 تل ole lot (38-4) y ((38-4) de + 2 to te - say i eny= 38€ - 4² y tn 38- 38 - J 2 Unis -ts iny

I use separation of variable to solve this problem